VIMOS News

On this page, we list the main changes that affected the instrument and this web site.

  • September 7, 2013: The report on slit losses as a function of setup and angle has been updated (Sanchez-Janssen et al. 2013). Basic conclusions remain the same: we accept E-W orientation for targets with declination between -45 and -5, for which HA can be +-3h from meridian. Caveat is that the acquisition should be performed in a filter whose wavelength is close to the grism wavelength used in science template.
  • February 12, 2013: mask2ccd and sky2ccd from May 2012 has been installed. Every pre-image starting tonight will have the new coefficients.
  • October 4, 2012: A modification of the focusing mechanism that allow a proper range for IFU focusing has been succesfully installed.
  • July 6, 2012: The VIMOS manual for P90 phase II has been released. It contains information on the new HR_blue VPH grisms, the new PILMOS mode, etc. You can retrieve it here.
  • June 11, 2012: A new Pre-Image Less MOS mode, or PILMOS mode, is being offered with a waiver during P90. See new manual.
  • June 11, 2012: The MOS mode restriction that slits must be oriented N-S has been somewhat relaxed. Now we accept E-W orientation for targets with declination between -45 and -5. HA can be +-3h from meridian. See short report by Sanchez-Janssen et al. (posted 2012-07-06). Note that this does not apply to the new PILMOS mode.
  • March 29, 2012: The focusing units were modified on March 15th to use motors with absolute encoders. At the same time the focusing software has been modified to be able to get optimal focus at all absolute rotator angles, and to adapt the focus to the grism used. The initial good news of image quality at the 0.45 arcsecond level in all quadrants over most of the field have been somewhat lessened by some instabilities detected in the active flexure compensation (AFC) system which share the same driver software as the new focusing units. We are currently investigating the causes. In the meantime operations are proceeding at a slower pace due to frequent failures.
  • March 29, 2012: A new set of HR_blue grisms has been installed in the instrument in March 15. The peak overall efficiency has been measured to be approximately 25%, compared with 15% previously (see MOS curves here). Preliminary information on the new HR-blue grisms characteristics (resolution, wavelength coverage, etc.) is available in the VIMOS User's Manual, and a more detailed report based on the observations will follow.
  • March 29, 2012: The mask cabinets have been upgraded to a new design of considerably reduced weight, redesigned mask holding magnet distribution, and a simpler installation mechanism. Although transparent to the users, we expect these modifications will increase the reliability of the instrument by improving mask positioning and reducing technical down time.
  • December 28, 2011: A new version of the VIMOS User manual has been made available here. This version contains preliminary information about the new VPH HR_blue grisms that will be used during this period.
  • June 19, 2011: New vmmps.cfg configuration file to be used in the post May 24 intervention improves a bit the area where slits can be placed. This file will be part of the new version of VMMPS (version 3.25) to be released in July 2011. The current version (version 3.24) can also beused for MOS OB preparation.
  • May 24, 2011: During the past dark time a new intervention on VIMOS improved significantly the focus of the instrument. In particular the ellipticity towards the corners has been reduced as can be seen from health check pages.In addition the active flexure compensation system is now operating during the entire OB.
  • Mar 24, 2011: New intervention of videoboard of FIERA-A (Q2,Q3). New gains and bias levels. Data taken after this date can not becalibrated with earlier data.
  • Mar 01, 2011: During twilight flats video board ofFIERA-A (Q2, Q3) burned out. Had to replace by spare. Gains and biaslevels changed. Notice that data taken after this date can not becalibrated with frames from earlier dates.
  • Feb 20, 2011: Bias levels of all quadrants adjusted from ~2400 ADU to ~200 ADU in all quadrants.
  • Feb 1, 2011: A VIMOS images among ESO's Hidden Treasures winners: http://www.eso.org/public/news/eso1102/
  • Oct 10, 2010: On the night of 2/3 Oct 2010 we have finished the verification tests that were needed during realignment work carried out the previous week. The instrument has been back in operations since then. Our priority since then has been to reduce the pre-imaging back log accumulated so far. All instrument modes are now operational. Currently ActiveFlexure Compensation is done only during acquisition, so we discourage the request of waivers for OBs lasting longer than 1h.
  • September 16, 2010: During September 21-27 the M1 ofUT3 will be recoated. During this time one further intervention to add guiding pins in Q2 will be done. This is done to ensure repeatability of cryostat mounting. We have postponed further pre-imaging until after the neccessary recharacterization after this intervention.
  • September 12, 2010: Focus has been checked and redetermined for all filters. New reference fiber position has beendefined for Q2. mask2ccd coefficients are in the process of being determined.
  • September 8, 2010: The Q2 cryostat is back into operations at the telescope.
  • September 6, 2010: The cryostat from Q2 has been taken out of operation and warmed up. This is done to correct an out of plane tilt of the CCD which is strongly affecting the focus in all filters.During this time we will only carry out observations in MOS HR and MRsettings.
  • August 30, 2010: Quadrants Q1 and Q2 have been realigned reducing the amount of CCD rotation. This intervention was needed to avoid the mismatch of the positions of the second orders ontothe first orders in LR multiplexed observations. These changes neccesitated new determinations of: (1) focus in all filters; (2) redetermination of AFC reference pixel; (3) redetermination of newmask2ccd coefficients; (4) redetermination of new WCS reference pixels;and (5) redetermination of new sky2ccd coefficients. No observations ofLR MOS OBs will be carried out until we can validate the above settings.
  • August 07, 2010: VIMOS is back on sky. The new CCDs have been installed and are in the process of being characterized on sky. Preliminary analysis indicate that efficiency has gone up by a factor of two for wavelengths longer than 800nm. Fringing in LR_red grism is not detectable, while for the HR_redgrism is approximately 2% peak-to-valey with a 1" slit, and 3% peak-to-valey with a 0.6" slit. The new Active Flexure Compensation mechanism is able to deliver corrections better than 0.3 pixels immediately after closing the loop. The open loop corrections are not implemented yet. During the month of August we will be doing pre-imaging and obtaining all calibration files needed by the new pipeline to be released soon. The new user manual should also be ready during this month. Stay tuned.
  • June 24, 2010: VIMOS is out of operations between May22 and July 31 due to the VIMOS Improvement Project. Afterwards VIMOSwill be equipped with new, more red sensitive detectors (see below) and an Active Flexure Compensation;
  • March 3, 2010: VIMOS Calibration Plan: P86 users are requested to attach to every MOS science OB a night calibration template with at least 3 flats and 1 arc;
  • September 4, 2009: VIMOS Improvement Project: preliminary QE curves ofthe new detectors can be found at http://www.eso.org/sci/facilities/paranal/instruments/vimos/inst/CCD_upgrade.html
  • September 4, 2009: Filter transmission curves and bandpasses for z-band have been added: filters.
  • March 26, 2009: VIMOS Improvement Project: We are currently defining an improvement project for VIMOS aimed principally to increase its red sensitivity by changing the detectors to e2v CCD44-82-1-D42, deep depletion, backside illuminated, double layer coating chips. We are also working on improving operation reliability. These changes will in principle be implemented during Q2 2010. Note that the blue sensitivity might decrease.
  • March 1, 2009: Starting P84 VIMOS User Manual, VMMPS Manual, and Guidecam Tool Manual are merged into a single document..
  • April 11, 2007: Instrument flexures have been measured and are reported in the web page: flexures.
  • April 11, 2007: Latest measurements of filter transmission curves and bandpasses: filters.
  • October 5, 2005: A new set of HR Red holographic grisms has been installed in VIMOS
  • September 1, 2005: These web pages have been updated
  • March 23, 2005: Slightly updated version of the P75 VMMPS Cookbook
  • March 21, 2005: Updated version of the User Manual for P76 CfP
  • February 15, 2005: release of the new web pages.
  • January 5, 2005: Updated version of the User Manual for P75 service
  • January 5, 2005: Updated version of the Guidecam User Manual for P75
  • June 20, 2004: Lines identification plots available